OpenStudy (princesssleelee):

David was looking at his limousine bill and saw that the total number of miles he has ridden in the car was tracked: Month Miles 5 580 6 695 7 810 8 925 The function representing David's miles per month is f(m) = 115m + 5. What does the 115 represent?

OpenStudy (princesssleelee):

The number of miles at the start The number of the month when David was riding The number of miles David had in total The number of additional miles each month

OpenStudy (princesssleelee):

im thinking its the starting miles, im not sure though

OpenStudy (texaschic101):

the number of the month for example : month 5.....sub in 5 for m f(5) = 115(5) + 5 = 575 + 5 = 580

OpenStudy (princesssleelee):

I understand that

OpenStudy (fortytherapper):

Correct @princesssleelee

OpenStudy (princesssleelee):

@texaschic101 do you agree?

OpenStudy (fortytherapper):

You stated the function was miles per month, which means f(m) equals miles

OpenStudy (fortytherapper):

The m represents the month and the 115 represents what he started with